OpenStudy (anonymous):

can someone please help me factor a polynomial that starts with a negative number? -2x^2-20x-48

OpenStudy (anonymous):

(-2x + )(x + ) or maybe (2x + )( -x + )

OpenStudy (phi):

factor out -2 from each term -2(x^2+10x+24) now factor x^2+10x+24

OpenStudy (anonymous):

Thanks Phi!! i was sooo close, but i just couldn't see it

OpenStudy (anonymous):

does the -2 just go away? or do I have to add it to the answer? ie -2(x^2-10-24) or does it just go away

OpenStudy (anonymous):

as usual, I am good, but Phi is better

OpenStudy (anonymous):

thanks anyway philips

OpenStudy (anonymous):

nothing just goes away, unless it's adding 0 or multiplying by 1 so keep the -2 as a factor

OpenStudy (anonymous):

ok thanks! I'm not so good at math yet... but with ya'lls help i'll get there!

OpenStudy (anonymous):

all it is, is practice... that is true of everything... the more you do, the better you get, at anything

OpenStudy (phi):

you get -2(x+4)(x+6) Now, if you had an equation -2x^2-20x-48=0 you would have -2(x+4)(x+6)=0 and you could divide both sides by -2 to get: (x+4)(x+6)=0 (In that sense, you can "make it go away") But when you just factor it, you keep the -2 out front.

OpenStudy (anonymous):

-(2x + 8)(x + 6) is the answer given and I can't figure out how to get there... any insight?

OpenStudy (anonymous):

i got what you got and then got a bit frustrated again...

OpenStudy (phi):

they just distributed the 2 over (x+4). in other words multiplied 2 times (x+4) they left the -1 out side; so -1(2x+8)(x+6)

OpenStudy (phi):

though you write it as -(2x+8)(x+6)

OpenStudy (phi):

Btw, my answer is "more factored" than the one given.

OpenStudy (anonymous):

can you show me the work from -2(x^2-4x=6x-24) to the answer given? i'm missing something... lol i have no doubts in your ability i'm still trying to grasp the basics... and they are eluding me still

OpenStudy (anonymous):

oops sorry i messed up writing the equation to you... sorry...

OpenStudy (phi):

-2x^2-20x-48 -2(x^2+10x+24) factor out -2 now factor the quadratic. The way I do it is: sign on 24 is + that means I want two numbers with the same sign. sign on 10 is +, so both numbers are + list the factors of 24: 1,24 no; 2,12 no; 3,8 no; 4,6 yes 4+6=10 so we get (x+4)(x+6) for (x^2+10x+24) remember the -2 to get -2(x+4)(x+6) if it is multiple choice and you see -(2x+8)(x+6) look at your answer and notice that -2(x+4)(x+6)= -(2x+8) (x+6)

OpenStudy (anonymous):

so initially when you factor out the -2 that would change the signs inside the parenthesis to =... that could be what i was missing

OpenStudy (phi):

yes, pulling out a minus sign from each term. You can check by multiplying by -2 and if you did it correctly, you get the original expression.

OpenStudy (anonymous):

good grief thank you so much... i think i might have it finally!! let the celebration commence